# Refusal Tokens: A Simple Way to Calibrate Refusals in Large Language Models

### v1

@@ Abstract
A key component of building safe and reliable language models is enabling the models to appropriately refuse to follow certain instructions or answer certain questions. We may want models to output refusal messages for various categories of user queries, for example, ill-posed questions, instructions for committing illegal acts, or queries which require information beyond the model’s knowledge horizon. Engineering models that refuse to answer such questions is complicated by the fact that an individual may want their model to exhibit varying levels of sensitivity for refusing queries of various categories, and different users may want different refusal rates. The current default approach involves training multiple models with varying proportions of refusal messages from each category to achieve the desired refusal rates, which is computationally expensive and may require training a new model to accommodate each user’s desired preference over refusal rates. To address these challenges, we propose refusal tokens, one such token for each refusal category or a single refusal token, which are prepended to the model’s responses during training. We then show how to increase or decrease the probability of generating the refusal token for each category during inference to steer the model’s refusal behavior. Refusal tokens enable controlling a single model’s refusal rates without the need of any further fine-tuning, but only by selectively intervening during generation.00footnotetext: Work completed during Capital One Internship; Correspondence to $<$njain17@umd.edu$>$
@@ 1 Introduction
An essential property of a useful language model is the ability to produce refusal messages at appropriate times. Refusal messages enhance not only the safety of LLMs but also their utility and trustworthiness, as refusal messages can prevent LLMs from hallucinating or answering invalid requests. For example, an LLM that lacks the ability to browse the web should refuse when asked to access and summarize the content behind a URL. Likewise, a model should provide an informative refusal when asked to answer a question that is too under-specified or poorly formed to be answerable. To minimize hallucinations and unsafe behavior, instruction models like GPT-4 (Achiam et al., 2023) and llama-3 (Dubey et al., 2024) have been processed with alignment pipelines that imbue them with extensive refusal capabilities. Despite advancements in model finetuning and alignment, controlling refusal messages in these models remains a challenging task. For instance, llama-2-Chat (Touvron et al., 2023) experienced issues with over-refusal, where the model would refuse too many queries, negatively impacting usability, mostly likely due to a post-training set with too many refusal messages. Simple approaches, such as training multiple models with varying levels of refusal data until the desired rates are achieved (Dubey et al., 2024) are resource-intensive and still lack the precision to carefully adjust different categories of refusals. Moreover, the criteria for refusing is constantly evolving. What is considered an acceptable refusal for one use case or time may not align with the ethical, legal, or technical standards in a different setting.
To address these weaknesses, we introduce a simple strategy that makes refusal behavior controllable at test-time without retraining: the refusal token. During alignment, we prepend a special [refuse] token to responses that contain a refusal. The model quickly learns to generate this token before refusing, and then to refuse when this token is present. At test-time, the softmax probability of the refusal token can be used as a metric for how likely it is that a refusal is necessary. By thresholding on this probability, one can turn a knob to control the refusal sensitivity after the model is trained. By employing different refusal tokens for different refusal types, one can impose fine-grained control over refusal behavior along different axes of behavior, and carefully optimize refusal rates in this multi-dimensional space.
@@
  • We introduce a refusal token strategy. By thresholding the probability of this refusal token, we give model developers calibrated control over refusal rates without retraining. This development opens the door for sophisticated post-training calibration of refusal rates. For example, with minimal computation, one could sweep over refusal thresholds and select a value that achieves a specified rate of false refusals, or a value that maximizes an F1 score.
  • We show that multiple refusal tokens can manage different refusal message sets, enabling independent control over each refusal distribution. Additionally, we explore various strategies for manipulating these category-specific refusal tokens to meet test-time requirements.
  • We observe that the refusal token improves F1 scores, even without calibration. Furthermore, we highlight the importance of reducing Type II errors by including contrast or borderline examples in the training data. These examples, which are similar to refusal queries but innocuous, help refine the token’s effectiveness—specifically, its ability to appropriately switch between refusal and response based on the corresponding meta-token.
@@ 2 Related Work
Refusal messages. The ability of generative models to refuse certain messages is particularly crucial for mitigating toxicity and reducing hallucinations. In the context of toxicity, several studies explore how language models respond to toxic prompts or instructions. One popular approach is to train an external model to determine whether the model should reject or respond to queries (Dubey et al., 2024). Bianchi et al. (2024) demonstrate that incorporating refusals into training data does not diminish a model’s helpfulness but can lead to over-refusals, where the model declines to respond even on innocuous requests. Similarly, Cui et al. (2024); An et al. (2024) investigate over-refusal behavior across various language models, developing an evaluation framework to assess over-refusals in response to harmful prompts. Regarding hallucinations, Zhang et al. (2024) introduce an algorithm called R-Tuning, which prompts the model to state “I am unsure” or “I am sure” after a question and answer session, framing the problem as a discrimination task. Additionally, Kang et al. (2024) and Kapoor et al. (2024) propose alternative algorithms for alleviating the hallucination problem, focusing on instances where it is unclear whether the model possesses the required knowledge. Feng et al. (2024) use multiple agents to determine when to abstain from queries. For predetermined queries the model is designed to refuse, Brahman et al. (2024) present a comprehensive taxonomy of such questions, highlighting scenarios where the model should appropriately refuse to respond. This work also releases instructional data designed to train models in this regard. Evaluative studies by Liu et al. (2023), Yin et al. (2023), and Amayuelas et al. (2024) further explore the types of questions that warrant refusal. Arditi et al. (2024) find a one-dimensional subspace such that erasing this specific direction from the model’s residual stream activations causes the model to consistently answer harmful queries. Concurrent work, Lee et al. (2024) extend Arditi et al. (2024) for usability, allowing additional steerability using activation steering to control refusal messages at test-time.
@@
Tagging, control codes, and meta-tokens. Sennrich et al. (2016) introduce a meta-token for machine translation, while Keskar et al. (2019) extend this idea by introducing control codes for more general control.. A control code is a piece of text, $c$, used in a conditional language model that always conditions on a control code $c$ and learns the distribution $p(x|c)$. Specifically, Keskar et al. (2019) pretrain a model using control codes to regulate style, content, and task-specific behavior. Tagging and control codes can also be viewed as form of prefix-tuning (Li & Liang, 2021). Dong et al. (2023) extend this idea by adding controls to different distributions during supervised fine-tuning (SFT) that users might want to control, including seven categories which are collected by training another classifier to first categorize and score the responses based on the selected seven attributes. These tags or tokens can also be predicted by the model to help the model generate its response to a query. The general use of these “meta-tokens”, or tokens that the model predicts to help itself generate its response to the query, has seen a recent increase with the introduction of tool calling in LLMs, or function calling (Nakano et al., 2021; Schick et al., 2024). However, others propose using meta-tokens for various purposes, such as enhancing reasoning capabilities (Yao et al., 2023), thinking capabilities (Goyal et al., 2024), or a variety of others (Teknium et al., 2024). In Table 1, we highlight the differences between these methods and our own.
@@ 3 Learning to Refuse with Tokens
Table 1: A list of capability differences between approaches applied to the language model for controlling refusal behavior. Refusal tokens provide more capabilities than other solutions. Tagging or control codes apply “tags” to the prompt to encourage safe outputs. In model reflection, the model outputs a response and then is asked to reflect on the safety of its response. Concurrent work introduced using activation steering (Lee et al., 2024) to control the refusal messages. See Section 2. Our proposed approach yields the more control over refusals: It (i) enables test-time control of the kinds of refusals that are enabled. It also (ii) produces an interpretable score (the refusal token “probability”) that quantifies the risk of answering without a refusal, and (iii) these scores can be thresholded/calibrated at inference time to optimize refusal rates. (iv) It also enables different refusal types/reasons to be adjusted separately. (v) It notifies the user with a special token when a refusal takes place, allowing developers to see the type query. Additionally, training with refusal tokens can improve F1 Scores without further calibrating the refusal token (i.e training time benefits).
User Input (Contains False Premise)
When did George Orwell write “The Invisible Man”?
Response (Low Refusal Threshold)
Response (High Refusal Threshold)
[respond] George Orwell wrote “The Invisible Man” in 1952.
Instruction models are trained on instruction-response pairs, $(x,y)$, sampled from an instruction dataset $D$. The user provides the model with a question or an instruction, $x$, and the model then outputs a response $y$. Each datapoint is usually given an additional chat template, $C$. Here, $y$ consists only of natural language without any meta-information contained in the messages. We introduce a new token, [refuse], at the beginning of the response if it is a refusal message, or [respond] otherwise during training. This modifies $y$ to $y^{\prime}=\texttt{[refuse]}+y$ or $y^{\prime}=\texttt{[respond]}+y$, depending on whether $y$ is a refusal message or a normal response.
We will see that including the [refuse] and [respond] tokens during training will influence the model at test-time. The model builds stronger associations during fine-tuning the more it encounters response tokens together with non-refusal messages and refusal tokens together with refusal messages. After fine tuning, the presence of the refusal token at the beginning of the response results in a high likelihood of a refusal message, and visa-versa.
@@
Test-time control. The primary reason to include this refusal token is the test-time capabilities that the token introduces. The model predicts this token, and there is a softmax probability associated with it that can be used as a confidence measure for determining whether the question should be refused or not. This confidence can manipulated in many ways such as thresholding the token or adding a logit bias. We focus our studies on the thresholding method, and emit the [refuse] token if its softmax score is $>T$, for some $T\in[0,1]$ chosen by the user.
Controlling different types of queries. We consider applying categorical refusal tokens for different refusal reasons. Our experimental setting includes five refusal tokens corresponding to the refusal categories defined in Brahman et al. (2024), and one respond token. Details of our multi-category thresholding schemes and logit bias mechanisms are described in greater detail in Section 5.1.
@@ 4 Experimental Set-up
We use the hyperparameters and codebase from Tunstall et al. (2023) for supervised finetuning. Our initial results with DPO (Rafailov et al., 2023) show that the SFT stage is required for the desired refusal behavior (See Appendix Table 6), and thus, we focus on the SFT stage for our experiments. The importance of the SFT stage before DPO for learning behaviors was studied in Sharma et al. (2024). We adopt llama-3 $8$B (Dubey et al., 2024) as the base model. Additionally, we mix the instruction pairs that contain refusal messages with UltraChat (Tunstall et al., 2023) or Alpaca (Taori et al., 2023). We experimented with Alpaca as it is largely free of any refusal messages, and its low training time facilitates more ablations in Section 6.
@@
CoCoNot Experimental Setting. For the main experimental setting, we utilize a diverse and comprehensive dataset—extending beyond just toxicity–for both training and evaluation to ensure robust performance in refusal prediction. Specifically, we adopt Brahman et al. (2024)’s CoCoNot dataset and evaluation due to the breadth of the categories and subcategories that are considered. The CoCoNot dataset contains five refusal categories–Humanizing, Indeterminate, Incomplete, Safety, and Unsupported–and $26$ subcategories. Additionally, the dataset contains contrast data, or queries/instructions that the model should answer but are close to questions that the model should refuse. We consider two main training settings on UltraChat with refusal data and training on UltraChat with refusal and contrast data. For these two settings, we either train with no refusal token, one refusal token, or multiple category refusal tokens. The CoCoNot dataset contains $\sim 10$k refusals SFT data, $\sim 1$k of contrast preference data (which we use as SFT data), and $\sim 1.4$k, or $1379$, for the evaluation. The evaluation contains $1$k queries that should refuse to answer and $379$ queries that the model should respond to the query–referred to as the contrast category. We refer to this evaluation and experimental set-up as CoCoNot.
[…]
@@
We consider two main training settings on UltraChat with refusal data and contrast data (Sections 5 and 6), and Alpaca (Taori et al., 2023) with refusal data and contrast data (Section 6). For these two settings, we either train with no refusal token or one refusal token. We consider this setting to understand the effect of balanced contrast data on the refusal token. For evaluation, we create $200$ temporal questions, which humans manually verified. In addition, the evaluation includes refusal instructions from CoCoNot’s refusal categories (excluding the temporal subcategory) and TriviaQA questions for model-appropriate responses. The inclusion of CoCoNot’s refusal questions is to determine how models may “generalize” to other refusal categories when trained only on a single question type (Section 6). The total query count is $1400$ for this evaluation, matching CoCoNot’s evaluation set. We refer to this evaluation and experimental set-up as Temp.
@@
Evaluation. For both experimental settings, we use the Brahman et al. (2024)’s prompts and evaluation framework with llama-3.1 $70$B as the LLM judge (Zheng et al., 2024). Brahman et al. (2024) originally found no quality difference between GPT-4 (Achiam et al., 2023) and GPT-3.5 (Brown, 2020). Furthermore, with llama-3.1 $70$B showing similar performance as GPT-3.5, we decided that an open-source model would be easier to reproduce as API models change and deprecate constantly. Additionally, we manually verified the effectiveness of llama-3.1 $70$B as the evaluator. Furthermore, we report F1 scores to three decimal places, as the standard error was measured to be below $0.002$ after both generation and evaluation.

### v2

@@ Abstract
A key component of building safe and reliable language models is enabling the models to appropriately refuse to follow certain instructions or answer certain questions. We may want models to output refusal messages for various categories of user queries, for example, ill-posed questions, instructions for committing illegal acts, or queries that require information beyond the model’s knowledge horizon. Engineering models that refuse to answer such questions is complicated by the fact that an individual may want their model to exhibit varying levels of sensitivity for refusing queries of various categories, and different users may want different refusal rates. The current default approach involves training multiple models with varying proportions of refusal messages from each category to achieve the desired refusal rates, which is computationally expensive and may require training a new model to accommodate each user’s desired preference over refusal rates. To address these challenges, we propose refusal tokens, one such token for each refusal category or a single refusal token, which are prepended to the model’s responses during training. We then show how to increase or decrease the probability of generating the refusal token for each category during inference to steer the model’s refusal behavior. Refusal tokens enable controlling a single model’s refusal rates without the need of any further fine-tuning, but only by selectively intervening during generation. Code is located at: github.com/neelsjain/refusal-tokens. 00footnotetext: Majority of work completed during Capital One Internship; Correspondence to <<njain17@umd.edu>>
@@ 1 Introduction
An essential property of a useful language model is the ability to produce refusal messages at appropriate times. Refusal messages enhance not only the safety of LLMs but also their utility and trustworthiness, as refusal messages can prevent LLMs from hallucinating or answering invalid requests. For example, an LLM that lacks the ability to browse the web should refuse when asked to access and summarize the content behind a URL. Likewise, a model should provide an informative refusal when asked to answer a question that is too under-specified or poorly formed to be answerable. To minimize hallucinations and unsafe behavior, instruction models like GPT-4 (Achiam et al., 2023) and llama-3 (Dubey et al., 2024) have been imbued with extensive refusal capabilities. Despite advancements in model finetuning and alignment, controlling refusal messages in these models remains a challenging task. For instance, llama-2-chat (Touvron et al., 2023) experienced issues with over-refusal, where the model would refuse too many queries, negatively impacting usability, mostly likely due to a post-training dataset with too many refusal messages. Simple approaches, such as training multiple models with varying levels of refusal data until the desired rates are achieved (Dubey et al., 2024) are resource-intensive and still lack the precision to carefully adjust different categories of refusals. Moreover, the criteria for refusing are constantly evolving. What is considered an acceptable refusal for one use case or time may not align with the ethical, legal, or technical standards in a different setting.
To address these weaknesses, we introduce a simple strategy that makes refusal behavior controllable at test-time without retraining: the refusal token. During alignment, we prepend a special [refuse] token to responses that contain a refusal. The model quickly learns to generate this token before refusing, and then to refuse when this token is present. At test-time, the softmax probability of the refusal token can be used as a metric for how likely it is that a refusal is necessary. By thresholding on this probability, one can turn a knob to control the refusal sensitivity without retraining. By employing different refusal tokens for different refusal types, one can impose fine-grained control over refusal behavior along different axes of behavior, and carefully optimize refusal rates in this multi-dimensional space.
@@
  • We introduce a refusal token strategy. By thresholding the probability of this refusal token, we give model developers calibrated control over refusal rates without retraining. This development opens the door for sophisticated post-training calibration of refusal rates. For example, with minimal computation, one could sweep over refusal thresholds and select a value that achieves a specified rate of false refusals, or a value that maximizes an F1 score. Alternately, an LLM user can adjust the refusal rate up or down just by “turning a knob.”
  • We show that multiple refusal tokens can manage different refusal message sets, enabling independent control over each refusal distribution. Additionally, we manipulate these category-specific refusal tokens to meet test-time requirements.
  • We observe that training with a refusal token improves F1 scores of refusal, even without calibration. Furthermore, we highlight the importance of reducing Type II errors by including contrast or borderline examples in the training data. These examples, which are similar to refusal queries but innocuous, help refine the token’s effectiveness—specifically, its ability to appropriately switch between refusal and response based on the corresponding meta-token.
@@ 2 Related Work
Refusal messages. The ability of generative models to refuse certain messages is particularly crucial for mitigating toxicity and reducing hallucinations. In the context of toxicity, several studies explore how language models respond to toxic prompts or instructions. One popular approach is to train an external model to determine whether the model should reject or respond to queries (Dubey et al., 2024). Bianchi et al. (2024) demonstrate that incorporating refusals into training data does not diminish a model’s helpfulness but can lead to over-refusals, where the model declines to respond even on innocuous requests. Similarly, Cui et al. (2024); An et al. (2024) investigate over-refusal behavior across various language models, developing an evaluation framework to assess over-refusals in response to harmful prompts. Regarding hallucinations, Zhang et al. (2024) introduce an algorithm called R-Tuning, which prompts the model to state “I am unsure” or “I am sure” after a question and answer session, framing the problem as a discrimination task. Additionally, Kang et al. (2024) and Kapoor et al. (2024) propose alternative algorithms for alleviating the hallucination problem, focusing on instances where it is unclear whether the model possesses the required knowledge. Feng et al. (2024) use multiple agents to determine when to abstain from queries. For predetermined queries the model is designed to refuse, Brahman et al. (2024) present a comprehensive taxonomy of such questions, highlighting scenarios where the model should appropriately refuse to respond. This work also releases instructional data designed to train models in this regard. Evaluative studies by Liu et al. (2023), Yin et al. (2023), and Amayuelas et al. (2024) further explore the types of questions that warrant refusal. Arditi et al. (2024) find a one-dimensional subspace such that erasing this specific direction from the model’s residual stream activations causes the model to consistently answer harmful queries. Concurrent work, Lee et al. (2024) extend Arditi et al. (2024) for usability, allowing additional steerability using activation steering to control refusal messages at test-time. Wen et al. (2025) present a comprehensive survey of refusal messages.
User Input (Contains False Premise)
When did George Orwell write “The Invisible Man”?
Response (Low Refusal Threshold)
Response (High Refusal Threshold)
[respond] George Orwell wrote “The Invisible Man” in 1952.
@@
Tagging, control codes, and meta-tokens. Sennrich et al. (2016) introduce a meta-token for machine translation, while Keskar et al. (2019) extend this idea by introducing control codes for more general control. A control code is a piece of text, $c$, used in a conditional language model that always conditions on a control code $c$ and learns the distribution $p(x|c)$. Specifically, Keskar et al. (2019) pretrain a model using control codes to regulate style, content, and task-specific behavior. Tagging and control codes can also be viewed as a form of prefix-tuning (Li & Liang, 2021). Dong et al. (2023) extend this idea by adding controls to different distributions during supervised fine-tuning (SFT) that users might want to control, including seven categories which are collected by training another classifier to first categorize and score the responses based on the selected seven attributes. These tags or tokens can also be predicted by the model to help the model generate its response to a query. The general use of these “meta-tokens”, or tokens that the model predicts to help itself generate its response to the query, has seen a recent increase with the introduction of tool calling in LLMs, or function calling (Nakano et al., 2021; Schick et al., 2024). However, others propose using meta-tokens for various purposes, such as enhancing reasoning capabilities (Yao et al., 2023), thinking capabilities (Goyal et al., 2024), or a variety of others (Teknium et al., 2024).
Table 1 highlights the differences between these methods and our own.
@@ 3 Learning to Refuse with Tokens
Table 1: A list of capability differences between approaches applied to the language model for controlling refusal behavior. Refusal tokens provide more capabilities than other solutions. Tagging or control codes apply “tags” to the prompt to encourage safe outputs. In model reflection, the model outputs a response and then is asked to reflect on the safety of its response. Concurrent work introduced using activation steering (Lee et al., 2024) to control the refusal messages. See Section 2. Our proposed approach yields more control over refusals: It (i) enables test-time control of the kinds of refusals that are enabled. It also (ii) produces an interpretable score (the refusal token “probability”) that quantifies the risk of answering without a refusal, and (iii) these scores can be thresholded/calibrated at inference time to optimize refusal rates. (iv) It also enables different refusal types/reasons to be adjusted separately. (v) It notifies the user with a special token when a refusal takes place, allowing developers to see the type of query. Additionally, training with refusal tokens can improve F1 Scores without further calibrating the refusal token (i.e, training time benefits).
Instruction models are trained on instruction-response pairs, $(x,y)$, sampled from an instruction dataset $D$. The user provides the model with a question or an instruction, $x$, and the model then outputs a response $y$. Each datapoint is usually given an additional chat template, $C$. Here, $y$ consists only of natural language without any meta-information contained in the messages. We introduce a new token, [refuse], at the beginning of the response if it is a refusal message, or [respond] otherwise during training. This modifies $y$ to $y^{\prime}=\texttt{[refuse]}+y$ or $y^{\prime}=\texttt{[respond]}+y$, depending on whether $y$ is a refusal message or a response faithfully answering the prompt.
We will see that including the [refuse] and [respond] tokens during training will influence the model at test-time. The model builds stronger associations during fine-tuning, the more it encounters response tokens together with non-refusal messages and refusal tokens together with refusal messages. After finetuning, the presence of the refusal token at the beginning of the response results in a high likelihood of a refusal message, and visa-versa.
@@
Test-time control. The primary reason to include this refusal token is the test-time capabilities that the token introduces. The model predicts this token, and thus, a softmax probability is associated with it that can be used as a confidence measure for determining whether the question should be refused or not. This confidence can be manipulated in many ways, such as thresholding the token or adding a logit bias. We focus our studies on the thresholding method, and emit the [refuse] token if its softmax score is greater than $T$, for some $T\in[0,1]$ chosen by the user.
Controlling different types of queries. We consider applying categorical refusal tokens for different refusal reasons. Our experimental setting includes five refusal tokens corresponding to the refusal categories defined in Brahman et al. (2024), and one respond token. Details of our multi-category thresholding scheme mechanism is described in greater detail in Section 5.1.
@@ 4 Experimental Set-up
We use the hyperparameters and codebase from Tunstall et al. (2023) for supervised finetuning. Our initial results with DPO (Rafailov et al., 2023) show that the SFT stage is required for the desired refusal behavior (See Appendix Table 6), and thus, we focus on the SFT stage for our experiments. The importance of the SFT stage before DPO for learning behaviors was studied in Sharma et al. (2024). We adopt llama-3 88B (Dubey et al., 2024) as the base model. Additionally, we mix the instruction pairs that contain refusal messages with UltraChat (Tunstall et al., 2023) or Alpaca (Taori et al., 2023). We experimented with Alpaca as it is largely free of any refusal messages, and its faster training time facilitates more ablations for Section 6.
@@
CoCoNot Experimental Setting. For the main experimental setting, we utilize a diverse and comprehensive dataset, extending beyond just toxicity, for both training and evaluation to ensure robust performance in refusal prediction. Specifically, we adopt Brahman et al. (2024)’s CoCoNot dataset and evaluation due to the breadth of its five categories (Humanizing, Indeterminate, Incomplete, Safety, and Unsupported), which encompass 26 subcategories. Additionally, the dataset contains contrast data, or queries/instructions that the model should answer, but are close to questions that the model should refuse. We consider two main training settings: UltraChat with refusal data and UltraChat with refusal and contrast data. For these two settings, we either train with no refusal token, a refusal and respond token, or multiple category refusal tokens with a respond token. The CoCoNot dataset contains $\sim 10$k refusals SFT examples, $\sim 1$k of contrast preference examples, which we use as SFT examples, and $\sim 1.4$k, or $1379$, for the evaluation. The evaluation consists of 1,000 queries to which the model should refuse to respond, and 379 queries to which it should provide a response, referred to as the contrast category.
Temporal Experimental Setting. We consider a second, more controlled experimental setting. We create temporal refusal and contrast training data to address CoCoNot’s low contrast-to-refusal ratio, at one to ten. For this setting, we consider a refusal message, where the query is temporally ambiguous or relates to events beyond the model’s cutoff dates. Additionally, we consider contrast data, or examples close to a refusal query but answerable, as temporal questions that contain dates about an event within its training period. Using llama-3 7070B, we prompt the model to generate questions from news articles beyond its cutoff date for refusal data, and before the cutoff date of the model for contrast data. More details are in Appendix A.4. We generate $\sim 2\text{k}$ examples each for refusal and contrast datasets, focusing on temporal questions, resulting in $\sim 4\text{k}$ instruction-response pairs.
@@
For the temporal setting, we experiment training on UltraChat with refusal data and contrast data (Sections 5 and 6) and Alpaca (Taori et al., 2023) with refusal data and contrast data (Section 6). For these two settings, we either train with no refusal token or one refusal and one respond token. We explore this setting to understand the effect of balanced contrast data on the refusal token. For the evaluation, we create $200$ temporal questions, which humans manually verified. In addition, the evaluation includes refusal instructions from CoCoNot’s refusal categories (excluding the temporal subcategory) and TriviaQA questions (Joshi et al., 2017) for model-appropriate responses. The inclusion of CoCoNot’s refusal questions is to determine how models may “generalize” to other refusal categories when trained only on a single question type (Section 6). The total query count is $\sim 1.4$k for this evaluation, matching CoCoNot’s evaluation set.
@@
Evaluation. For both CoCoNot and the temporal experimental settings, we use the Brahman et al. (2024)’s prompts and evaluation framework with llama-3.1 7070B as the LLM judge (Zheng et al., 2024). Brahman et al. (2024) report no quality difference between GPT-4 (Achiam et al., 2023) and GPT-3.5 (Brown, 2020). Furthermore, with llama-3.1 7070B showing similar performance as GPT-3.5 (White et al., 2025), we decided that an open-source model would be easier to reproduce as API models change and deprecate constantly. Additionally, we manually verify the effectiveness of llama-3.1 7070B as the evaluator. Following the rubric provided by CoCoNot, we label 150 randomly sampled examples from the llama-3 + UltraChat baseline. We find that CoCoNot achieved approximately $91\%$ agreement, and the temporal evaluation achieves approximately $95\%$ agreement. Additionally, we find from the $13$ incorrect annotations for CoCoNot, all but one of them is a qualified answer marked as respond when the rubric points that the label should be a refusal. Furthermore, we report F1 scores to three decimal places, as the standard error is measured to be below $0.002$ after both generation and evaluation.